# Open-Sora 2.0: Training a Commercial-Level Video Generation Model in \$200k

### v2

@@ 2.3 Data Statistics
3 Model Architecture
3.1 3D Autoencoder
Model Down. (TxHxW) Info. Down. Channel Causal LPIPS↓ PSNR↑ SSIM↑
Open Sora 1.2 [zheng2024open] $4\times 8\times 8$ 192 4 $\checkmark$ 0.161 27.504 0.756
StepVideo VAE [ma2025step] $8\times 16\times 16$ 96 64 $\checkmark$ 0.082 28.719 0.818
HunyuanVideo VAE [kong2024hunyuanvideo] $4\times 8\times 8$ 48 16 $\checkmark$ 0.046 30.240 0.856
Video DC-AE $4\times 32\times 32$ 96 128 $\times$ 0.051 30.538 0.863
Video DC-AE $4\times 32\times 32$ 48 256 $\times$ 0.049 30.777 0.872
We begin by leveraging the open-source HunyuanVideo VAE [kong2024hunyuanvideo] as the initial autoencoder for our model. To further reduce both training and inference costs, we develop a video autoencoder with deep compression (Video DC-AE, named after [chen2024deep]) that improves efficiency while maintaining high reconstruction fidelity.
HunyuanVideo VAE achieves a compression ratio of $4\times 8\times 8$, which allows an 8-second, high-resolution ($1280\times 720$) video at 16 FPS to be transformed into a latent representation of 32×160×90. However, even with a patch size of $2\times 2$, our generation model still needs to process around 115K tokens per training video, resulting in considerable computational demands.
To address the above challenge, we identify a potential redundancy in the spatial dimension and propose to increase the spatial compression ratio to 32 while keeping the same temporal compression ratio at 4. This adjustment effectively reduces the number of spatial tokens while preserving essential motion features in our Video DC-AE. Since our AE training data consists of 32-frame, 256px videos, this new compression ratio allows us to train the autoencoder at a latent shape of $8\times 8\times 8$, which maintains the same level of information in all dimensions.
[…]
chen2024deep introduce DC-AE as an effective approach to achieve higher downsampling ratios while maintaining reconstruction quality. While the original DC-AE is primarily optimized for image encoding, we extend its architecture to support video encoding by incorporating temporal compression mechanisms.
As shown in Figure 5(a), our Video DC-AE encoder consists of three residual blocks followed by three EfficientViT blocks [cai2023efficientvit], with the decoder adopting a symmetrical structure. The first five encoder blocks and the last five decoder blocks function as downsampling and upsampling layers, respectively.
To adapt DC-AE for video encoding, we replace 2D operations (e.g., convolutions, normalization) with 3D operations. Additionally, we introduce temporal compression in the last two downsampling blocks (blocks 4 and 5) of the encoder, and temporal upsampling in the first two upsampling blocks (blocks 2 and 3) of the decoder to reconstruct temporal information effectively.
[…]
As shown in Figure 5(b), our downsample residual blocks follow DC-AE’s pixel-shuffling strategy, redistributing pixels from the spatial and temporal dimensions into the channel dimension, followed by channel-wise averaging to achieve the desired compression. The upsample residual blocks perform the inverse operation by first duplicating channels, and subsequently redistributing them back into the spatial and temporal dimensions to reconstruct the original structure.
We train Video DC-AE from scratch and evaluate its reconstruction quality, with results summarized in Table 1. Our model is compared against OpenSora 1.2 VAE, the recently open-sourced StepVideo VAE [ma2025step], and HunyuanVideo VAE [kong2024hunyuanvideo]. The results demonstrate that Video DC-AE achieves competitive performance, with only minor degradation in LPIPS [zhang2018unreasonable] scores compared to the best-performing HunyuanVideo VAE, while maintaining strong PSNR and SSIM scores.
3.2 DiT Architecture
Figure 6: Open-Sora diffusion transformer architecture.
To achieve higher video quality, we employ full attention to capture long-range dependencies effectively. After encoding the input through the autoencoder, we patchify the latent representations to enhance computational efficiency and improve model learning. Sana [xie2024sana] suggests that a patch size of 1 yields better training stability and finer details in video generation. With Video DC-AE’s high compression ratio, we can afford to reduce the patch size to 1 (i.e., no patch at all), whereas a patch size of 2 is still required when the HunyuanVideo autoencoder is used in our generation model.
Inspired by FLUX’s MMDiT [flux2024], we employ a hybrid transformer architecture that incorporates both dual-stream and single-stream processing blocks. In the dual-stream blocks, text and video information are processed separately to facilitate more effective feature extraction within each modality. Subsequently, single-stream blocks integrate these features to facilitate effective cross-modal interactions. To further enhance the model’s ability to capture spatial and temporal information, we apply 3D RoPE (Rotary Position Embedding) [su2024roformer], which extends traditional positional encoding to three-dimensional space, allowing the model to better represent motion dynamics across time.
For text encoding, we leverage T5-XXL [chung2024scaling] and CLIP-Large [Radford2021LearningTV], two high-capacity pretrained models known for their strong semantic understanding. T5-XXL captures complex textual semantics, while CLIP-Large improves alignment between text and visual concepts, leading to more accurate prompt adherence in video generation.
An overview of the model architecture is illustrated in Figure 6, and the detailed architectural specifications are presented in Table 2.
@@ 3.2 DiT Architecture
4 Model Training
A commercial-grade video generation model typically requires more than 10 billion parameters and O(100M) training samples. To mitigate the substantial computational cost associated with training, we propose a cost-effective training pipeline, as outlined in Table 3. Our approach consists of three key stages: (1) training a text-to-video model on low-resolution video data, (2) training an image-to-video model on low-resolution video data, and (3) fine-tuning an image-to-video model on high-resolution videos.
Training Stage Dataset CP #iters #GPUs #GPU day USD
256px T2V 70M 1 85k 224 2240 \$107.5k
256px T/I2V 10M 1 13k 192 384 \$18.4k
768px T/I2V 5M 4 13k 192 1536 \$73.7k
Total 4160 \$199.6k
4.1 Efficient Training Strategy
To develop a high-quality video generation model within a constrained budget, our training strategy emphasizes the following four key aspects:
Leveraging Open-Source Image Models
Prior research [ma2024latte, hong2022cogvideo, zheng2024open] has demonstrated that pretraining on image datasets can significantly accelerate video model training. To avoid the high cost of training an image model from scratch, we leverage an open-source solution. Specifically, we adopt Flux [flux2024], a state-of-the-art text-to-image model with 11 billion parameters, which provides sufficient capacity to generate high-quality videos. We initialize our text-to-video model using Flux and empirically find that, despite it being a distilled model, this initialization is effective for further training.
High-Quality Training Data
Inspired by the efficient image training strategies from PixArt [chen2023pixart], we hypothesize that high-quality video data can substantially enhance training efficiency. Consequently, we curate a high-quality subset from a large-scale dataset for low-resolution training. For high-resolution fine-tuning, we impose stricter selection criteria to ensure superior video quality.
Learning Motion in Low-Resolution
As shown in Table 4 and Table 5, training on a 129-frame video at 768px resolution is 40 times slower than at 256px. This performance gap arises due to the quadratic computational complexity of self-attention mechanisms as the number of tokens increases. To optimize efficiency, we allocate the majority of computational resources to low-resolution training, reducing the need for expensive high-resolution computations.
Image-to-Video Models Facilitate Resolution Adaptation
We find that adapting a model from 256px to 768px resolution is significantly more efficient using an image-to-video approach compared to text-to-video. We hypothesize that conditioning the model on a static image allows it to focus more on motion generation, a capability that is well-learned during low-resolution training.
Based on this observation, we prioritize training an image-to-video model at high resolution. During inference, we first generate an image from a text prompt and subsequently synthesize a video conditioned on both the image and the text. During training, text/image-to-video training at low resolution followed by brief fine-tuning on high-resolution videos yields high-quality results with minimal additional training.
With our proposed approach, we successfully constrain the one-time training cost to \$200K, as detailed in Table 3. We estimate the training costs of Movie Gen [polyak2024moviegen] and Step-Video-T2V [ma2025step] based on publicly available information, as summarized in Figure 7. Our model achieves 5–10× lower training costs. We hope that these insights will contribute to reducing the training cost of high-quality video generation models in future research.
4.2 Training Setting
Our training setup is largely based on our previous iteration, Open-Sora 1.2 [zheng2024open]. We adopt flow matching [lipman2022flow] as our primary training objective and utilize the AdamW [loshchilov2017decoupled] optimizer with $\beta$ values of (0.9, 0.999) and an $\epsilon$ value of $1\times 10^{-15}$. No weight decay is applied. The learning rate is set to $5\times 10^{-5}$ for the first 40k steps of Stages 1 and 2, followed by a decay to $3\times 10^{-5}$ for the final 45k steps. For Stage 3, it is reduced to $1\times 10^{-5}$. Gradient norm clipping is applied with a threshold of 1.
Training Objective
Our flow matching approach is similar to that used in Stable Diffusion 3 [esser2024scaling]. We denote the video latent as $X_{0}$, and Gaussian noise $X_{1}\sim\mathcal{N}(0,1)$. The model $f_{\theta}$ takes as input an interpolated latent $X_{t}=(1-t)X_{0}+tX_{1}$ and is trained to predict the velocity component $X_{0}-X_{1}$. The corresponding loss function is formulated as:
where $y$ represents the conditioning input (text and/or image). The timestep $t$ is first sampled from a logit-normal distribution and then scaled according to the shape of $X_{0}$. Given that higher-resolution and longer-duration videos are more susceptible to noise, we apply the following transformation:
$t\leftarrow\frac{\alpha t}{1+(\alpha-1)t},$
where $\alpha$ is proportional to the product $T\times H\times W$. The same method is applied during inference.
Multi-Bucket Training
resolution #frames max #tokens batch size throughput on 8 GPUs
@@ 4.2 Training Setting
Table 4: Batch size and throughput in stage 1 and 2.
resolution #frames max #tokens batch size throughput on 8 GPUs
@@
Following the methodology in Open-Sora 1.2 [zheng2024open], we adopt multi-bucket training to efficiently handle videos of varying frame counts, resolutions, and aspect ratios within the same batch. This strategy optimizes GPU utilization by dynamically assigning batch sizes based on video characteristics. The specific batch sizes used for different training configurations are detailed in Table 4 and Table 5.
@@
4.3 High Compression Autoencoder Adaptation
Advantages.
Utilizing a high-compression video autoencoder significantly reduces the number of tokens required for video generation while increasing the number of latent channels incurs minimal additional computational cost (affecting only the input and output layers). We define the token downsample ratio, $D_{\text{token}}$, as the product of the autoencoder’s compression ratios ($D_{\text{T | H | W}}$) and the generation model’s patch sizes ($P_{\text{T | H | W}}$) in the T, H, W dimensions:
@@ 4.3 High Compression Autoencoder Adaptation
For instance, in a 5-second, 24fps, 768px video, the HunyuanVideo VAE with a $4\times 8\times 8$ compression in combination of generation model’s patch size $1\times 2\times 2$ ($D_{\text{token}}=1024$) reduces the token count to 76K, whereas the Video DC-AE with $4\times 32\times 32$ compression and patch size of $1\times 1\times 1$ has a $D_{\text{token}}$ of 4096, effectively reduces the token count to 19K. As shown in Figure 8, this compression yields a 5.2×5.2$\times$5.2 × speedup in training throughput. Additionally, inference speeds improve by over 10×10$\times$10 ×, making it a highly efficient approach.
Challenges in Training Video Autoencoders.
With the growing interest in video generation models, we observe that despite variations in architecture, downsampling ratios, number of channels, and training losses, video autoencoders trained with similar information downsampling ratios exhibit comparable performance (evaluated at 256px resolution) [chen2024deep]. We define the information downsampling ratio, $D_{\text{info}}$ as follows:
However, training a deep video compression network is challenging. Prior work [chen2024deep] identifies gradient propagation issues in downsample and upsample blocks. To address this, they introduce residual connections in these blocks using pixel shuffle and unshuffle operations. We verify that training the HunyuanVideo VAE architecture with residual connections improves performance. Building on this, we incorporate temporal downsampling into DC-AE to construct Video DC-AE.
Additionally, we find that maintaining similar reconstruction performance requires a 4× increase in channels when doubling the height and width compression ratio, whereas adding a 4× temporal compression on top of the original image DC-AE (no temporal compression) incurs no additional channels possibly due to redundancy in temporal information. Based on these insights, we train a Video DC-AE with a 2x $D_{\text{info}}$ and 16x $D_{\text{token}}$ as compared to the HunyuanVideo VAE while maintaining comparable performances.
Challenges in Training the Generation Model.
[…]
Strategies for Using High-Compression Autoencoders.
  1. 1.
    Latent Space Distillation: After training Video DC-AE, we apply a distillation loss to align the third-layer latents with DINOv2 [oquab2023dinov2], leveraging their similar latent shapes.
  2. 2.
  3. 3.
    Prioritizing Image-to-Video Training: We observe that image-to-video models adapt more efficiently than text-to-video models when switching to high-compression AEs. Thus, we focus on training an image-to-video model for this setting.
  4. 4.
[…]
We first train the model on short videos (up to 33 frames) for 17K iterations using 20M samples across 160 GPUs. Then, we extend training to long videos (up to 128 frames) for 8K iterations on 2M samples. The final model achieves a loss level of 0.5, compared to 0.1 for the initialization model. However, due to computational constraints, training does not fully converge. As shown in Figure 9, while the fast video generation model underperforms the original, it still captures spatial-temporal relationships. We release this model to the research community for further exploration.
Discussion and Future work.
We believe high-compression autoencoders are critical for end-to-end video generation, especially for high-resolution videos that contain substantial redundancy. To further reduce training costs, it is essential to: (1) Optimize autoencoder training to yield better latent spaces for diffusion-based learning. (2) Design high-throughput autoencoders, as autoencoder encoding time remains a major computational bottleneck in generation models using high-compression autoencoders.
4.4 AE Training
$\mathcal{L}=\mathcal{L}_{1}+0.5\mathcal{L}_{\text{LPIPS}},$
5 Conditioning
5.1 Image-to-Video Training
Figure 10: Image and Video Condition Framework Condition information is injected into the model via an additional channel dimension. A mask mechanism is introduced to distinguish different input types. This design enables support for a wide range of image-to-video (I2V) and video-to-video (V2V) tasks.
As illustrated in Figure 10, our approach first encodes image or video conditions using an autoencoder, then concatenates the encoded features with the original video latent representation. An extra channel is introduced to indicate the task type, increasing the number of channels from $k$ to $12k+$ 1 .
To improve generalization, we introduce image condition dropout similar to text condition dropout. During training, dropping the image condition reduces the problem to a text-to-video setting, where zero tensors are concatenated to the video latent. For T/I2V training, we set the dropout ratio to 12.5%, ensuring robustness across various image-to-video (I2V) and text-to-video (T2V) tasks.
5.2 Image-to-Video Inference
Figure 11: Heatmap of Image Guidance Scale Across Denoising Steps and Latent Frames Darker regions indicate higher image guidance values, emphasizing stronger influence on later frames and earlier denoising steps.
We use classifier-free guidance for inference [ho2022classifier]. As our model is conditioned on both image and text, a straightforward approach is to use a single guidance scale:
$v_{t}$ $=v_{\theta}(x_{t},t,\varnothing,\varnothing)$
A further issue arises when using high image guidance, as it sometimes introduces flickering in generated videos. To mitigate this, we introduce a guidance oscillation technique [ho2022imagen] for image guidance. Taking 50-step sampling as an example, after the first 10 steps, we alternate the image guidance scale: during odd-numbered steps, we apply the original $g_{\textit{img}}$, while for even-numbered steps, we reduce it to 1. This helps balance stability and motion consistency.
5.3 Motion Score
Controlling the level of motion strength is a crucial feature in video generation. Depending on the scene, users may prefer high-fidelity videos with minimal motion or highly dynamic videos with significant movement. While techniques such as classifier-free guidance scaling and shifting denoising steps can influence motion intensity, we find that these approaches are often entangled with other factors, such as visual quality and text alignment, making precise control challenging. To address this, we explicitly model the motion dynamics as a separate controllable parameter.
We leverage the motion score obtained from data pre-processing, which quantifies the level of dynamics in a video. This motion score is appended to the caption as an additional conditioning signal. During inference, adjusting the motion score allows for effective and independent control over the dynamic level of the generated video, as demonstrated in Figure 12.
5.4 Inference-Time Scaling
@@
[…]
These candidates are then evaluated using six VBench metrics [huang2024vbench]: subject_consistency, background_consistency, motion_smoothness, dynamic_degree, aesthetic_quality, imaging_quality. Depending on the application, users can assign higher priority to specific metrics—for example, favoring motion_smoothness for more natural movements or aesthetic_quality for improved visual appeal. The highest-scoring candidate is selected to continue in the denoising process, allowing the model to explore a richer generation space with stronger alignment to target attributes.
@@
While inference-time scaling improves overall video quality, we do not include it in our evaluation comparisons with other models to ensure a fair and consistent benchmarking setting.
@@
6 System Optimization
We train our models using ColossalAI [10.1145/3605573.3605613], an efficient parallel training system. Our hardware setup includes H200 GPUs, whose 141GB memory enables more effective data parallelism (DP) and allows for more aggressive selective activation checkpointing, significantly optimizing resource utilization. Additionally, we leverage PyTorch compile [Ansel_PyTorch_2_Faster_2024] and Triton kernels [triton_language] to accelerate training efficiency.
6.1 Parallelism Strategy
To efficiently handle high-resolution video training, we employ multiple parallelization techniques. For video autoencoders, we adapt tensor parallelism (TP) [shoeybi2020megatronlmtrainingmultibillionparameter] to convolution layers by partitioning weights either the input or output channel dimensions, reducing memory consumption while preventing out-of-bound indexing for high-resolution training.
For MMDiT training, we combine Zero Redundancy Optimizer (ZeroDP) [rajbhandari2020zeromemoryoptimizationstraining] with Context Parallelism (CP) [liu2023ringattentionblockwisetransformers]. Here, video and text sequences are partitioned across GPUs along the sequence dimension, enabling each GPU to compute attention independently. This approach effectively mitigates the memory bottleneck and enhances the efficiency of attention computation, which is particularly beneficial for high-resolution videos, where attention complexity grows quadratically.
Empirical evaluations on H200 GPUs (141GB memory) indicate that employing CP alone yields an optimal trade-off between memory efficiency and computational performance. During Stage 1 and Stage 2 training, we exclusively utilize data parallelism (DP) in combination with ZeRO-2, achieving a maximum FLOPs utilization (MFU) of 38.19%. In Stage 3, we integrate ZeRO-2 with CP=4, resulting in to an MFU of 35.75%.
@@ 6.1 Parallelism Strategy
6.2 Activation Checkpointing
6.3 Auto Recovery
6.4 Dataloader
6.5 Checkpoint Optimization
7 Performance
Our model supports both text-to-video and image-to-video generation at 256×256px and 768×768px resolutions (hereafter referred to as 256px and 768px), generating videos up to 128 frames long. At 24 FPS, this corresponds to a 5-second duration. Since our model is optimized for image-to-video generation, the default text-to-image-to-video pipeline first generates an image using the FLUX model, which is then used as the starting frame for video generation. The generated results are presented in Figure 15.
We further evaluate our model’s performance using VBench [huang2024vbench] in Figure 16, demonstrating significant improvements from Open-Sora 1.2 to 2.0. The performance gap between Open-Sora and OpenAI’s Sora has been reduced from 4.52% to 0.69%, highlighting substantial advancements in video generation quality. Additionally, our model achieves a higher VBench score compared to CogVideoX1.5-5B and HunyuanVideo, further establishing its superiority among current open-source text-to-video models.
Figure 16: VBench Score Comparison. Our model outperforms open-source text-to-video models by leveraging a text-to-image-to-video generation approach. Additionally, our latest version significantly narrows the performance gap between Open-Sora and OpenAI’s Sora, demonstrating substantial improvements in video generation quality and coherence.
8 Conclusion
This report introduces Open-Sora 2.0, a commercial-level video generation model that was trained for only \$200k, which is 5-10 times more cost-efficient than comparable models like MovieGen and Step-Video-T2V. This achievement highlights that high-quality video generation models can be developed with highly controlled costs through careful optimization of data curation, model architecture, training strategy, and system optimization. Despite its significantly lower training cost, Open-Sora 2.0 performs comparably to leading video generation models including HunyuanVideo and Runway Gen-3 Alpha. The model supports both text-to-video and image-to-video generation at resolutions up to 768×768 pixels for videos up to 5 seconds in length.
Looking ahead, several challenges remain in video generation. First, deep compression video VAE technology is still underexplored. While we aggressively increase the compression ratio to reduce latent tokens, this approach introduce reconstruction quality loss and adaptation difficulties. Additionally, diffusion models often produce unpredictable artifacts such as object distortion and unnatural physics, with users having limited control over these details. The field needs further research on artifact prevention and enhanced control over generated content. We hope that by open-sourcing Open-Sora 2.0, we can provide the community with tools to collectively tackle these challenges, fostering innovation and advancements in the field of video generation.
@@ 8 Conclusion
  • Contributors: Chaoyu Gong, Guojun Lei, Leijun Cheng, Limin Zhang, Minghao Li, Ruijie Zhang, Silan Hu, Shijie Huang, Xiaokang Wang, Yuanheng Zhao, Yuqi Wang, Ziang Wei.

### v3

@@ Appendix A Data Statistics
Appendix B Data Filters
B.1 Aesthetic Score.
We assess the aesthetic quality of a sample using the CLIP+MLP aesthetic score predictor (Schuhmann, 2021). For a video sample, we extract the first, middle, and last frame, compute their individual scores, and take the average as the final aesthetic score.
B.2 Motion Score
The motion intensity of a video is measured using the VMAF motion score from FFmpeg’s libavfilter library (FFmpeg Developers, 2023). Videos with extremely low or excessively high motion scores are filtered out to maintain a balanced motion range.
B.3 Blur Detection
To evaluate image clarity, we apply the Laplacian operator from OpenCV (Bradski, 2000). A sample is considered blurry if the variance of the Laplacian image falls below a configurable threshold. For video samples, we extract five uniformly spaced frames and adopt a majority voting approach to determine blurriness.
B.4 OCR
We detect text bounding boxes using PaddleOCR (Authors, 2020) and compute the total area of bounding boxes with a confidence score above 0.7. Images or videos containing excessive text are discarded to avoid unwanted overlays.
B.5 Camera Jitter Detection
We detect camera jitter using Shot Boundary Detection from the PySceneDetect library (Castellano, 2023). A video is identified as having camera jitter if the average frame-to-frame change exceeds a predefined threshold.
Appendix C Model Architecture
C.1 Architecture Hyperparameters
@@
Appendix D Multi-Bucket Training
For multi-bucket training, we adapt different batch sizes for different training configurations, such as parallelism strategy, video resolution, number of frames, and maximum number of frames. The specific batch sizes used for different training configurations are detailed in Table 4 and Table 5.
Table 4: Batch size and throughput in stage 1 and 2.
Resolution # Frames Max # of Frames Batch Size Throughput on 8 GPUs
@@ Appendix D Multi-Bucket Training
Resolution # Frames Max # of Frames Batch Size Throughput on 8 GPUs
@@
Appendix E Comparison of Video Generation with Different Autoencoder Compression Ratios
Appendix F High Compression Autoencoder Adaptation
F.1 Derivation of The Token Count Reduction Effects
[…]
@@ F.1 Derivation of The Token Count Reduction Effects
For instance, in a 5-second, 24fps, 768px video, the HunyuanVideo VAE with a $4\times 8\times 8$ compression in combination of generation model’s patch size $1\times 2\times 2$ ($D_{\text{token}}=1024$) reduces the token count to 76K, whereas the Video DC-AE with $4\times 32\times 32$ compression and patch size of $1\times 1\times 1$ has a $D_{\text{token}}$ of 4096, effectively reduces the token count to 19K.
Appendix G Inference-Time Scaling
@@
We explore an inference-time scaling strategy inspired by Ma et al. (2025b) to enhance video generation quality and diversity without modifying the underlying model. As illustrated in Figure 14, this approach introduces controlled noise variations during key denoising steps—particularly early steps (e.g., steps 1 and 3), which significantly influence the final output. At each selected step, partial denoising is performed to generate multiple candidate outputs, each corresponding to a different noise variant.
These candidates are then evaluated using six VBench metrics (Huang et al., 2024): subject_consistency, background_consistency, motion_smoothness, dynamic_degree, aesthetic_quality, imaging_quality. Depending on the application, users can assign higher priority to specific metrics—for example, favoring motion_smoothness for more natural movements or aesthetic_quality for improved visual appeal. The highest-scoring candidate is selected to continue in the denoising process, allowing the model to explore a richer generation space with stronger alignment to target attributes.
@@
Appendix H Other System Optimization
H.1 Activation Checkpointing
H.2 Auto Recovery
H.3 Dataloader
H.4 Checkpoint Optimization
Appendix I Generation Samples
@@
  • Contributors: Chaoyu Gong, Guojun Lei, Leijun Cheng, Limin Zhang, Minghao Li, Ruijie Zhang, Silan Hu, Shijie Huang, Xiaokang Wang, Yuanheng Zhao, Yuqi Wang, Yuxuan Lou, Ziang Wei.